knowing a rate of return of 6% and a standard deviation of 13%. What percentage of returns were greater than 45%
1
Expert's answer
2020-04-09T14:21:30-0400

Assume that returns of this portfolio "X" follow a Normal distribution: "X\sim N(\mu, \sigma^2)."

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given that "\mu=0.06, \sigma=0.13."


"P(X>0.45)=1-P(X\leq0.45)=1-P(Z\leq{0.45-0.06\over 0.13})=""=1-P(Z\leq3)\approx1-0.998650\approx0.001350"

"0.135\%"